# Score Forgetting Distillation: A Swift, Data-Free Method for Machine Unlearning in Diffusion Models

### v2

@@ B.1 Datasets for class forgetting tasks
For the class forgetting tasks, we utilize CIFAR-10 [Krizhevsky, 2009] at a resolution of $32\times 32$ and STL-10 [Coates et al., 2011] at $64\times 64$ resolution. The CIFAR-10 dataset consists of 60,000 32$\times$32 color images in 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images. The dataset consists of 50,000 training images and 10,000 test images. It is organized into five training batches and one test batch, each containing 10,000 images. The test batch includes precisely 1,000 randomly-selected images from each class. The training batches, which hold the remaining images in random order, may have varying numbers of images from each class. The STL-10 dataset is another natural image dataset with 10 classes, each of which has 500 training data and 800 testing data. The image data has a higher resolution of 96$\times$96 in pixels and RGB color channels compared with CIFAR-10. The images were acquired from labeled examples on ImageNet [Deng et al., 2009]. During training time, the image data from STL-10 are resized to 64$\times$64. Due to the limited number of the original training data, both training and testing data were used in the experiments, making up 13,000 training images in total.
@@ B.2 Evaluation
For class forgetting tasks, we employed an external classifier to obtain unlearning accuracy (UA), ensuring that the generated images are not associated with the class or concept designated for forgetting. The UA is essentially the mis-classification rate of the classifier on the generated samples from the target class. A classifier with high test accuracy and low UA typically indicates effective forgetting, ensuring that the generated images are unlikely to belong to the target class or concept. For the external classifier, we fine-tuned ResNet-34 [He et al., 2016] for 10 epochs on both CIFAR-10 and STL-10 datasets using transfer learning, which is originally pretrained on ImageNet [Deng et al., 2009]. We adapted the original 1000-way classification model by replacing the last fully-connected layer with a customized fully-connected layer with 10 output dimension. The resulting classifiers achieved training and testing accuracies of 99.96% and 95.03% on CIFAR-10, and 100.00% and 96.20% on STL-10, respectively.
@@
We followed the Inappropriate Image Prompts (I2P) benchmark introduced by Schramowski et al. [2023] to assess the risk of generating NSFW images in text-to-image diffusion models. The I2P dataset consists of 4,703 text prompts covering a wide range of NSFW concepts, including “nudity.” For each prompt, we generated 10 images and applied both the NudeNet and Q16 detectors to identify inappropriate content. We report the sample-level inappropriate probability (referred to as “Inapprop. Prob.”) and the prompt-level inappropriate rate (referred to as “Max. Exp. Inapprop.”).
@@ B.4 Implementation Details
We implemented our techniques in a newly developed codebase, loosely based on the original implementations by [Karras et al., 2022, Fan et al., 2024, Zhou et al., 2024a]. The pseudo-code is described in Algorithm 1. We performed extensive evaluation to verify that our implementation produced exactly the same results as previous work, including samplers, pre-trained models, network architectures, training configurations, and evaluation. We ran all experiments using PyTorch with 4 NVIDIA RTX A5000 GPUs.
@@ B.6 Forgetting Nudity as a Concept
B.7 Hyperparameter Settings
@@ Appendix D Proof of Lemma 1
$E_{g_{\theta}}\|s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2})\|^{2}$ $=E_{g_{\theta}}[(s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2}))^{T}s_{\phi}]-E_{g_{\theta}}[(s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2}))^{T}s_{\theta}]$

### v3

@@ B.1 Datasets for class forgetting tasks
For the class forgetting tasks, we utilize CIFAR-10 [41] at a resolution of $32\times 32$ and STL-10 [14] at $64\times 64$ resolution. The CIFAR-10 dataset consists of 60,000 32$\times$32 color images in 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images. The dataset consists of 50,000 training images and 10,000 test images. It is organized into five training batches and one test batch, each containing 10,000 images. The test batch includes precisely 1,000 randomly-selected images from each class. The training batches, which hold the remaining images in random order, may have varying numbers of images from each class. The STL-10 dataset is another natural image dataset with 10 classes, each of which has 500 training data and 800 testing data. The image data has a higher resolution of 96$\times$96 in pixels and RGB color channels compared with CIFAR-10. The images were acquired from labeled examples on ImageNet [15]. During training time, the image data from STL-10 are resized to 64$\times$64. Due to the limited number of the original training data, both training and testing data were used in the experiments, making up 13,000 training images in total.
@@ B.2 Evaluation
For class forgetting tasks, we employed an external classifier to obtain unlearning accuracy (UA), ensuring that the generated images are not associated with the class or concept designated for forgetting. The UA is essentially the mis-classification rate of the classifier on the generated samples from the target class. A classifier with high test accuracy and low UA typically indicates effective forgetting, ensuring that the generated images are unlikely to belong to the target class or concept. For the external classifier, we fine-tuned ResNet-34 [26] for 10 epochs on both CIFAR-10 and STL-10 datasets using transfer learning, which is originally pretrained on ImageNet [15]. We adapted the original 1000-way classification model by replacing the last fully-connected layer with a customized fully-connected layer with 10 output dimension. The resulting classifiers achieved training and testing accuracies of 99.96% and 95.03% on CIFAR-10, and 100.00% and 96.20% on STL-10, respectively.
@@
We followed the Inappropriate Image Prompts (I2P) benchmark introduced by [65] to assess the risk of generating NSFW images in text-to-image diffusion models. The I2P dataset consists of 4,703 text prompts covering a wide range of NSFW concepts, including “nudity.” For each prompt, we generated 10 images and applied both the NudeNet and Q16 detectors to identify inappropriate content. We report the sample-level inappropriate probability (referred to as “Inapprop. Prob.”) and the prompt-level inappropriate rate (referred to as “Max. Exp. Inapprop.”).
@@ B.4 Implementation Details
We implemented our techniques in a newly developed codebase, loosely based on the original implementations by [39; 19; 95]. The pseudo-code is described in Algorithm 1. We performed extensive evaluation to verify that our implementation produced exactly the same results as previous work, including samplers, pre-trained models, network architectures, training configurations, and evaluation. All experiments were conducted using four NVIDIA RTX A5000 GPUs. For class forgetting tasks, we pretrained base diffusion models for CIFAR-10 and STL-10. For concept forgetting tasks, including celebrity forgetting and nudity forgetting under the SFD-Two Stage setting, we utilized the pretrained checkpoint provided by Zhou et al. [95], which is a one-step diffusion model based on Stable Diffusion 1.5.
@@ B.6 Forgetting Nudity as a Concept
B.7 Additional Experiments on Adversarial Robustness
We conducted additional experiments to evaluate the robustness of our method against adversarial attacks. Specifically, we followed the adversarial setup described in UnlearnDiffAtk [91] and evaluated our nudity-forgetting SFD model under scenarios without attacks and with adversarial prompts. We measured the adversarial robustness of our model using the Attack Success Rate (ASR), calculated based on NudeNet detection results of generated images from 142 prompts in the I2P dataset. We denote the scenario without attacks as “Pre-ASR” and the scenario with UnlearnDiffAtk as “Post-ASR.”
In addition to MU baselines for diffusion models, we included a stronger baseline in terms of adversarial robustness against UnlearnDiffAtk, $i.e.,$ AdvUnlearn [90]. Here, “AdvUnlearn-UN” and “AdvUnlearn-TE” represent SD models with UNet and text encoder finetuned using AdvUnlearn, respectively. The evaluation results are provided in Table 8.
We note that our method, SFD, achieves the best Pre-ASR among all baselines and the best Post-ASR among all UNet-based baselines, demonstrating the inherent robustness of our model. While the original SFD model underperforms AdvUnlearn-TE in Post-ASR, incorporating AdvUnlearn-TE into our SFD model (referred to as “SFD-TE”) achieves the best adversarial robustness across all models. These results further demonstrate the flexibility and adaptability of our method.
Table 8: Adversarial robustness of different MU methods
Metric ESD FMN SLD AdvUnlearn-UN AdvUnlearn-TE SFD (ours) SFD+TE (ours) Pre-ASR 20.42% 88.03% 33.10% - 7.75% 7.04% 0.70% Post-ASR 76.05% 97.89% 82.39% 64.79% 21.13% 55.63% 7.04%
B.8 Additional Experiments on Alternative Score Distillation Methods
To demonstrate the flexibility of the SFD framework, we adapted it to accommodate alternative score distillation methods, such as Diff-Instruct [48]. Specifically, we incorporated a Kullback-Leibler (KL)-divergence-based forgetting distillation loss into the SFD framework, resulting in a variant denoted as “SFD-KL.” Similar to Equation 3, this KL-based score forgetting distillation loss is defined as follows:
$\mathcal{L}_{\text{sfd-kl}}(\theta;\phi,c_{1},c_{2})=\mathbb{E}_{z_{t},t,x\sim\mathcal{D}_{\theta,c_{2}}}\left[\omega_{t}\log\frac{p_{\theta}(z_{t}\mid c_{2},t)}{p_{\phi}(z_{t}\mid c_{1},t)}\right].$ (13)
Following Luo et al. [48], the gradient of this loss with respect to the generator is given by:
$\nabla_{\theta}\mathcal{L}_{\text{sfd-kl}}=\mathbb{E}_{z_{t},t,x\sim\mathcal{D}_{\theta,c_{2}}}\left[\omega_{t}\alpha_{t}\left(s_{\psi^{*}(\theta)}(z_{t},c_{2},t)-s_{\phi}(z_{t},c_{1},t)\right)\nabla_{\theta}x\right],$ (14)
where the true score $s_{\psi^{*}}$ is approximated by the score estimator $s_{\psi}$ during training. A comparison of the original SFD and the adapted SFD-KL is presented in Table 9. While both methods perform well, the original SFD achieves superior results across all metrics except Precision, demonstrating its enhanced generation quality and diversity. Figure 9 further highlights the advantages of SFD, showcasing faster convergence and a lower final FID compared to SFD-KL. These findings emphasize the efficiency and effectiveness of SFD in MU tasks. Overall, the results demonstrate the flexibility of the SFD framework in adapting to alternative score distillation methods while maintaining competitive performance. Additionally, the faster convergence and improved generation quality achieved by the original SFD underscore its robustness and practicality for real-world applications.
Table 9: Comparison of SFD and SFD-KL across various metrics.
Model UA (↑) FID (↓) IS (↑) Precision (↑) Recall (↑) NFEs (↓) Data-free
SFD 99.64 5.35 9.51 0.6587 0.5471 1 Yes
SFD-KL 99.53 6.99 9.44 0.6688 0.5016 1 Yes
Figure 9: FID training curve comparison between SFD and SFD-KL. SFD achieves faster convergence and a better final FID, highlighting its superior efficiency and performance.
B.9 Hyperparameter Settings
@@ Appendix D Proof of Lemma 1
$E_{g_{\theta}}\|s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2})\|^{2}$
$=E_{g_{\theta}}[(s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2}))^{T}s_{\phi}]-E_{g_{\theta}}[(s_{\phi}(y,c_{1})-s_{\theta}(y,c_{2}))^{T}s_{\theta}]$